# Generalizable Spectral Embedding with an Application to UMAP

### v1

@@ Abstract
Spectral Embedding (SE) is a popular method for dimensionality reduction, applicable across diverse domains. Nevertheless, its current implementations face three prominent drawbacks which curtail its broader applicability: generalizability (i.e., out-of-sample extension), scalability, and eigenvectors separation. In this paper, we introduce GrEASE: Generalizable and Efficient Approximate Spectral Embedding, a novel deep-learning approach designed to address these limitations. GrEASE incorporates an efficient post-processing step to achieve eigenvectors separation, while ensuring both generalizability and scalability, allowing for the computation of the Laplacian’s eigenvectors on unseen data. This method expands the applicability of SE to a wider range of tasks and can enhance its performance in existing applications. We empirically demonstrate GrEASE’s ability to consistently approximate and generalize SE, while ensuring scalability. Additionally, we show how GrEASE can be leveraged to enhance existing methods. Specifically, we focus on UMAP, a leading visualization technique, and introduce NUMAP, a generalizable version of UMAP powered by GrEASE. Our codes are publicly available.  [^1] GrEASE: https://github.com/shaham-lab/GrEASE; NUMAP: https://github.com/shaham-lab/NUMAP
@@ 1 Introduction
Spectral Embedding (SE) is a popular non-linear dimensionality reduction method (Belkin and Niyogi, 2003; Coifman and Lafon, 2006a), finding extensive utilization across diverse domains in recent literature. Notable applications include UMAP (McInnes et al., 2018) (the current state-of-the-art visualization method), Graph Neural Networks (GNNs) (Zhang et al., 2021; Beaini et al., 2021) and Graph Convolutional Neural Networks (GCNs) (Defferrard et al., 2016), positional encoding for Graph Transformers (Dwivedi and Bresson, 2020; Kreuzer et al., 2021) and analysis of proteins (Campbell et al., 2015; Kundu et al., 2004; Shepherd et al., 2007; Zhu and Schlick, 2021). The core of SE involves a projection of the samples into the space spanned by the leading eigenvectors of the Laplacian matrix (i.e., those corresponding to the smallest eigenvalues), derived from the pairwise similarities between the samples. It is an expressive method which is able to preserve the global structure of high-dimensional input data, underpinned by robust mathematical foundations (Belkin and Niyogi, 2003; Katz et al., 2019; Lederman and Talmon, 2018; Ortega et al., 2018).
@@
Despite the popularity and significance of SE, current implementations suffer from three main drawbacks: (1) Generalizability - the ability to directly embed a new set of test points after completing the computation on a training set (i.e., out-of-sample extension); (2) Scalability - the ability to handle a large number of samples within a reasonable timeframe; (3) Eigenvectors separation - the ability to output the basis of the leading eigenvectors $(v_{2},\dots,v_{k+1})$, rather than only the space spanned by them. These three properties are crucial for modern applications of SE in machine learning. Notably, the last property has attracted considerable attention in recent years (Pfau et al., 2018; Gemp et al., 2020; Deng et al., 2022; Lim et al., 2022). While most SE implementations address two of these three limitations, they often fall short in addressing the remaining one (see Tab. 1 and Sec. 2).
@@
This paper extends the work by Shaham et al. (2018), known as SpectralNet. SpectralNet tackles the scalability and generalizability limitations of Spectral Clustering (SC), a key application of SE. However, we prove that due to a rotation and reflection ambiguity in its loss function, SpectralNet cannot directly be adapted for SE (i.e., it cannot separate the eigenvectors). In this paper, we first present a post-processing procedure to resolve the eigenvectors separation issue in SpectralNet, thereby, creating a scalable and generalizable implementation of SE, which we call GrEASE: Generalizable and Efficient Approximate Spectral Embedding.
GrEASE’s ability to separate the eigenvectors, while maintaining generalizability and scalability offers a pathway to enhance numerous existing applications of SE and provides a foundation for developing new applications. A notable example is UMAP (McInnes et al., 2018), the current state-of-the-art visualization method. A recent work by Sainburg et al. (2021) proposed Parametric UMAP (P. UMAP) to address UMAP’s lack of generalizability. However, UMAP’s global structure preservation and consistency largely stem from the use of SE for initialization (Kobak and Linderman, 2021), a step absent in P. UMAP. Consequently, P. UMAP lacks a crucial component to fully replicate the performance of UMAP, especially in terms of global structure preservation. Nonetheless, a series of studies have incorporated P. UMAP, underscoring the significant impact of a generalizable version of UMAP (Xu and Zhang, 2023; Eckelt et al., 2023; Leon-Medina et al., 2021; Xie et al., 2023; Yoo et al., 2022).
In this paper, we also introduce a novel application of GrEASE for generalizable UMAP, which we term NUMAP. NUMAP integrates the UMAP loss with SE initialization, similar to the original non-parametric UMAP. As a result, NUMAP achieves comparable results to UMAP, while also offering generalization capabilities. This extends UMAP applicability, for instance, to the online learning regime and visualization of time-series.
Our contributions can be summarized as follows: (1) We introduce GrEASE, a novel approach for generalizable approximate SE; (2) We establish a foundation for a range of new SE applications and enhancements to existing methods; (3) We present NUMAP: a novel application of GrEASE for generalizable UMAP; (4) We propose a new evaluation method for dimensionality reduction methods, which enables quantification of global structure preservation.
@@ 2 Related Work
Table 1: GrEASE is the only method to have the three desired properties of SE implementation. Comparison between key SE methods via their ability to generalize to unsee samples, scale to large datasets and separate the eigenvectors.
@@
GrEASE (ours)
@@
Popular implementations of SE are mostly based on sparse matrix decomposition techniques (e.g., ARPACK (Lehoucq et al., 1998), AMG (Brandt et al., 1984), LOBPCG (Benner and Mach, 2011)). These methods are relatively scalable, as they are almost linear in the number of samples. Nevertheless, their out-of-sample extension is far from trivial. Usually, it is done by out-of-sample extension (OOSE) methods such as Nyström (Nyström, 1930) or Geometric Harmonics (Coifman and Lafon, 2006b; Lafon et al., 2006). However, these methods provide only local extension (i.e., near existing training points), and are both computationally and memory restrictive, as they rely on computing the distances between every new test point and all training points.
@@
Several approaches to SC approximate the space spanned by the first eigenvectors of the Laplacian matrix, which is sufficient for clustering purposes, and can also benefit other specific applications. For example, SpectralNet (Shaham et al., 2018) leverages deep neural networks to approximate the first eigenfunctions of the Laplace-Beltrami operator in a scalable manner, thus also enabling fast inference of new unseen samples. BASiS (Streicher et al., 2023) achieves these goals using affine registration techniques to align batches. However, these methods’ inability to separate the eigenvectors prevents their use in many modern applications.
@@
Another proposed approach to SE is DiffusionNet (Mishne et al., 2019), a deep-learning framework for generalizable Diffusion Maps embedding (Coifman and Lafon, 2006a), which is similar to SE. However, the training procedure of the network is computationally expensive, therefore restricting its usage for large datasets.
In contrast, we introduce GrEASE, which generalizes the separated eigenvectors to unseen points with a single feed-forward operation, while maintaining scalability.
@@
Extensive research has been conducted on the eigenvectors separation problem, both within and beyond the spectral domain (Lim et al., 2022; Ma et al., 2024). However, recent suggestions are constrained computationally, both by extensive run-time and memory consumption. For example, Pfau et al. (2018) proposed a solution to this issue by masking the gradient information from the loss function. However, this approach necessitates the computation of full Jacobians at each time step, which is highly computationally intensive. Gemp et al. (2020) employs an iterative method to learn each eigenvector sequentially. Namely, they learn an eigenvector while keeping the others frozen. This process has to be repeated $k$ times (where $k$ is the embedding dimension), which makes this approach also computationally expensive. Deng et al. (2022) proposed an improvement of the latter, by parallel training of $k$ NNs. However, as discussed in their paper, this approach becomes costly for large values of $k$. Furthermore, it necessitates retaining $k$ trained networks in memory, which leads to significant memory consumption. Chen et al. (2022) proposed a post-processing solution to this problem using the Rayleigh-Ritz method. However, this approach involves the storage and multiplication of very large dense matrices, rendering it impractical for large datasets. In contrast, GrEASE offers an efficient one-shot post-processing solution to the eigenvectors separation problem.
@@
Several works have attempted to develop parametric approximations for non-parametric visualization methods, in addition to Parametric UMAP (P. UMAP) (Sainburg et al., 2021). Notable examples include (Van Der Maaten, 2009) and (Kawase et al., 2022), which use NNs to make t-SNE generalizable, and (Schofield and Lensen, 2021), which aims to make UMAP more interpretable. However, P. UMAP has demonstrated superior performance. NUMAP presents a method to surpass P. UMAP in terms of global structure preservation.
@@ 3.1 Spectral Embedding
Let $\mathcal{X}=\{x_{1},\dots,x_{n}\}\subseteq\mathbb{R}^{d}$ denote a collection of unlabeled data points drawn from some unknown distribution $\mathcal{D}$. Let $W\in\mathbb{R}^{n\times n}$ be a positive symmetric graph affinity matrix, with nodes corresponding to $\mathcal{X}$, and let $D$ be the corresponding diagonal degree matrix (i.e. $D_{ii}=\sum_{j=1}^{n}W_{ij}$). The Unnormalized Graph Laplacian is defined as $L=D-W$. Other normalized Laplacian versions are the Symmetric Laplacian $L_{\text{sym}}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ and the Random-Walk (RW) Laplacian $L_{\text{rw}}=D^{-1}L$. GrEASE is applicable to all of these Laplacian versions. The eigenvalues of $L$ can be sorted to satisfy $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ with corresponding eigenvectors $v_{1},\dots,v_{n}$ (Von Luxburg, 2007). It is important to note that the first pair (i.e., $\lambda_{1},v_{1}$) is trivial - for every Laplacian matrix $\lambda_{1}=0$, and for the unnormalized and RW Laplacians $v_{1}=\frac{1}{\sqrt{n}}\vec{1}$, namely the constant vector.
@@ 3.2 SpectralNet
The Rayleigh quotient (RQ) of a Laplacian matrix $L\in\mathbb{R}^{n\times n}$ is a function $R_{L}:\mathbb{R}^{n\times k}\to\mathbb{R}$ defined by
$R_{L}(A)=\text{Tr}(A^{T}LA)$
@@
$\mathcal{L}_{\text{spectralnet}}(\theta)=\frac{1}{m^{2}}R_{L}\big{(}f_{\theta}(X)\big{)}\quad s.t.\quad\frac{1}{m}f_{\theta}(X)^{T}f_{\theta}(X)=I_{k\times k}$ (1)
Thereby, it learns a map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ (where $d$ is the input dimension) which approximates the space spanned by the first $k$ eigenfunctions of the Laplace-Beltrami operator on the underlying manifold $\mathcal{D}$ (Belkin and Niyogi, 2006; Shi, 2015). Following this, it clusters the representations via KMeans. These eigenfunctions are a natural generalization of the SE to unseen points, enabling both scalable and generalizable spectral clustering.
@@ 3.3 UMAP and Parametric UMAP
Figure 1: NUMAP preserves global structure while enabling generalizability. (a) A comparison between non-parametric UMAP (with SE or PCA initialization), P. UMAP and NUMAP on three non-linear yet simple 3-dimensional toy datasets. NUMAP global structure abilities over P. UMAP are evident. (b) Banknote’s test set visualization by P. UMAP, NUMAP-SN and NUMAP. A better separation between classes is observed in NUMAP.
UMAP (McInnes et al., 2018) is the current state-of-the-art visualization method. It presented a significant advancement over previous methods, primarily due to its enhanced scalability and superior ability to preserve global structure. This approach involves the construction of a graph from the input high-dimensional data and the learning of a low-dimensional representation. The objective is to minimize the KL-divergence between the input data graph and the representations graph.
However, as discussed in (Kobak and Linderman, 2021), UMAP primarily derives its global preservation abilities, as well as its consistency, from initializing the representations using SE. Therefore, the SE initialization serves as a critical step for UMAP to uphold the global structure (see demonstration in Fig. 1a). Global preservation, in this context, refers to the separation of different classes, and avoiding the separation of existing classes. We refer the reader to (Kobak and Linderman, 2021) for a more comprehensive discussion about the effects of informative initialization on UMAP’s performance.
UMAP method can be divided into three components (summarized in Fig. 3): (1) constructing a graph which best captures the global structure of the input data; (2) initializing the representations via SE; (3) Learning the representations, via SGD, which best capture the original graph. This setup does not facilitate generalization, as both steps (2) and (3) lack generalizability.
Recently, a generalizable version of UMAP, known as Parametric UMAP (P. UMAP), was introduced (Sainburg et al., 2021). P. UMAP replaces step (3) with the training of a neural network. Importantly, it overlooks step (2), the SE initialization. Consequently, P. UMAP struggles to preserve global structure, particularly when dealing with non-linear structures. Fig. 1a illustrates this phenomenon with several non-linear yet simple structures. Noticeably, P. UMAP fails to preserve global structure (e.g., it does not separate different clusters).

### v2

@@ Abstract
Spectral Embedding (SE) is a popular method for dimensionality reduction, applicable across diverse domains. Nevertheless, its current implementations face three prominent drawbacks which curtail its broader applicability: generalizability (i.e., out-of-sample extension), scalability, and eigenvectors separation. Existing SE implementations often address two of these drawbacks; however, they fall short in addressing the remaining one. In this paper, we introduce Sep-SpectralNet (eigenvector-separated SpectralNet), a SE implementation designed to address all three limitations. Sep-SpectralNet extends SpectralNet with an efficient post-processing step to achieve eigenvectors separation, while ensuring both generalizability and scalability. This method expands the applicability of SE to a wider range of tasks and can enhance its performance in existing applications. We empirically demonstrate Sep-SpectralNet’s ability to consistently approximate and generalize SE, while maintaining SpectralNet’s scalability. Additionally, we show how Sep-SpectralNet can be leveraged to enable generalizable UMAP visualization. [^1] Sep-SpectralNet: https://github.com/shaham-lab/GrEASE; NUMAP: https://github.com/shaham-lab/NUMAP
@@ 1 Introduction
Spectral Embedding (SE) is a popular non-linear dimensionality reduction method (Belkin & Niyogi, 2003; Coifman & Lafon, 2006b), finding extensive utilization across diverse domains in recent literature. Notable applications include UMAP (McInnes et al., 2018) (the current state-of-the-art visualization method), Graph Neural Networks (Zhang et al., 2021; Beaini et al., 2021; Defferrard et al., 2016), signal propagation on graphs (Park et al., 2022; Klaine et al., 2017) and analysis of proteins (Campbell et al., 2015; Zhu & Schlick, 2021). The core of SE involves a projection of the samples into the space spanned by the leading eigenvectors of the Laplacian matrix (i.e., those corresponding to the smallest eigenvalues), derived from the pairwise similarities between the samples. It is an expressive method which is able to preserve the global structure of high-dimensional input data, underpinned by robust mathematical foundations (Belkin & Niyogi, 2003; Katz et al., 2019; Lederman & Talmon, 2018; Ortega et al., 2018).
@@
Despite the popularity and significance of SE, current implementations suffer from three main drawbacks: (1) Generalizability - the ability to directly embed a new set of test points after completing the computation on a training set (i.e., out-of-sample extension); (2) Scalability - the ability to handle a large number of samples within a reasonable time-frame; (3) Eigenvectors separation - the ability to output the basis of the leading eigenvectors $(v_{2},\dots,v_{k+1})$, rather than only the space spanned by them. These three properties are crucial for modern applications of SE in machine learning. While most SE implementations address two of these three limitations, they often fall short in addressing the remaining one (see Tab. 1 and Sec. 2).
Notably, eigenvector separation has attracted considerable attention in recent years (Pfau et al., 2018; Gemp et al., 2020; Deng et al., 2022; Lim et al., 2022). This work focuses on this property, which is a key aspect underlying applications such as Fiedler vector, Diffusion maps, and Diffusion analysis. The latter two involve examining the evolution of random walks on graphs across different time scales. This framework enables a wide range of applications: analyzing molecular transitions (Glielmo et al., 2021; Chiavazzo et al., 2017), sampling rare protein transitions (Ghamari et al., 2024), uncovering latent structures in disordered materials (Hardin et al., 2024), studying the global organization of cortical features in various neuropsychiatric conditions (Park et al., 2021; 2022; Dong et al., 2020), supporting enhanced-sampling techniques (Zheng et al., 2011; 2013), and facilitating dimensionality reduction in self-organizing networks (Klaine et al., 2017). Eigenvector separation is required for those applications to enable scaling each diffusion component.
@@
Shaham et al. (2018) presented SpectralNet, which tackles the scalability and generalizability limitations of Spectral Clustering (SC), a key application of SE. However, we prove that due to a rotation and reflection ambiguity in its loss function, SpectralNet cannot directly be adapted for SE in general, as it cannot separate the eigenvectors. In this paper, we present an eigendecomposition based post-processing procedure to resolve the eigenvectors separation issue in SpectralNet, thereby, extending SpectralNet into a scalable and generalizable implementation of SE, which we call Sep-SpectralNet (eigenvector-separated SpectralNet). Sep-SpectralNet’s ability to separate the eigenvectors, while maintaining the generalizability and scalability of SpectralNet, offers a pathway to enhance numerous existing applications of SE, that require eigenvector separation, and provides a foundation for developing new applications.
One such application is UMAP (McInnes et al., 2018), a widely used visualization method whose ability to preserve global structure relies heavily on its Spectral Embedding (SE) initialization (Kobak & Linderman, 2021). Although Parametric UMAP (P. UMAP) was proposed to address UMAP’s lack of generalizability (Sainburg et al., 2021), it omits SE, limiting its ability to replicate UMAP’s structural fidelity. Despite this limitation, P. UMAP has gained traction in various domains (Xu & Zhang, 2023; Eckelt et al., 2023; Leon-Medina et al., 2021; Xie et al., 2024; Yoo et al., 2022), underscoring the need for a generalizable alternative that retains UMAP’s structural strengths. To this end, we introduce an extension of Sep-SpectralNet that integrates SE-based initialization with the UMAP loss, named NUMAP, preserving global structure while enabling generalization. We show that eigenvector separation is required for NUMAP’s performance. As shown empirically (Sec. 5), this extension enhances UMAP’s applicability to dynamic settings such as online learning and time-series data visualization.
Our key contributions are: (1) We introduce Sep-SpectralNet, a novel approach for approximate Spectral Embedding (SE) that jointly addresses scalability, generalizability, and eigenvector separation. (2) We present an application of Sep-SpectralNetto generalizable UMAP. (3) We propose a new evaluation method to measure global structure preservation.
@@ 2 Related Work
Sep-SpectralNet (ours)
Table 1: Sep-SpectralNet is the only method to have the three desired properties of SE implementation. Comparison between key SE methods via their ability to generalize to unseen samples, scale to large datasets and separate the eigenvectors.
@@
Popular implementations of SE are mostly based on sparse matrix decomposition techniques (e.g., ARPACK (Lehoucq et al., 1998), AMG (Brandt et al., 1984), LOBPCG (Benner & Mach, 2011)). These methods are relatively scalable, as they are almost linear in the number of samples. Nevertheless, their out-of-sample extension is far from trivial. Usually, it is done by out-of-sample extension (OOSE) methods such as Nyström (Nyström, 1930) or Geometric Harmonics (Coifman & Lafon, 2006a; Lafon et al., 2006). However, these methods provide only local extension (i.e., near existing training points), and are both computationally and memory restrictive, as they rely on computing the distances between every new test point and all training points.
@@
Several approaches to spectral clustering (SC) approximate the space spanned by the first eigenvectors of the Laplacian matrix, which is sufficient for clustering purposes, and can also benefit other specific applications. For example, SpectralNet (Shaham et al., 2018) leverages deep neural networks to approximate the first eigenfunctions of the Laplace-Beltrami operator in a scalable manner, thus also enabling fast inference of new unseen samples. BASiS (Streicher et al., 2023) achieves these goals using affine registration techniques to align batches. However, these methods’ inability to separate the eigenvectors prevents their use in many modern applications.
@@
Another proposed approach to SE is DiffusionNet (Mishne et al., 2019), a deep-learning framework for generalizable Diffusion Maps embedding (Coifman & Lafon, 2006b), which is similar to SE. However, the training procedure of the network is computationally expensive, therefore restricting its usage for large datasets.
In contrast, we introduce Sep-SpectralNet, which addresses all three limitations - generalizes the separated eigenvectors to unseen points with a single feed-forward operation, while maintaining SpectralNet’s scalability.
@@
Extensive research has been conducted on the eigenvectors separation problem, both within and beyond the spectral domain (Lim et al., 2022; Ma et al., 2024). Some rotation criteria such as ICA and VARIMAX are well known, but regarding the spectral domain, they do not yield the natural separation, i.e., the true eigenvectors. Recent spectral approaches remain constrained computationally, both by extensive run-time and memory consumption. For example, Pfau et al. (2018) proposed a solution to this issue by masking the gradient information from the loss function. However, this approach necessitates the computation of full Jacobians at each time step, which is highly computationally intensive. Gemp et al. (2020) employs an iterative method to learn each eigenvector sequentially. Namely, they learn an eigenvector while keeping the others frozen. This process has to be repeated $k$ times (where $k$ is the embedding dimension), which makes this approach also computationally expensive. Deng et al. (2022) proposed an improvement of the latter, by parallel training of $k$ NNs. However, as discussed in their paper, this approach becomes costly for large values of $k$. Furthermore, it necessitates retaining $k$ trained networks in memory, which leads to significant memory consumption. Chen et al. (2022) proposed a post-processing solution to this problem using the Rayleigh-Ritz method. However, this approach involves the storage and multiplication of very large dense matrices, rendering it impractical for large datasets. In contrast, Sep-SpectralNet offers an efficient one-shot post-processing solution to the eigenvectors separation problem.
@@
Several works have attempted to develop parametric approximations for non-parametric visualization methods, in addition to Parametric UMAP (P. UMAP) (Sainburg et al., 2021). Notable examples include (Van Der Maaten, 2009), (Kawase et al., 2022) and (Damrich et al., 2022), which use NNs to make t-SNE generalizable, and (Schofield & Lensen, 2021), which aims to make UMAP more interpretable. However, P. UMAP has demonstrated superior performance. NUMAP presents a method to surpass P. UMAP in terms of global structure preservation.
@@ 3.1 Spectral Embedding
Let $\mathcal{X}=\{x_{1},\dots,x_{n}\}\subseteq\mathbb{R}^{d}$ denote a collection of unlabeled data points drawn from some unknown distribution $\mathcal{D}$. Let $W\in\mathbb{R}^{n\times n}$ be a positive symmetric graph affinity matrix, with nodes corresponding to $\mathcal{X}$, and let $D$ be the corresponding diagonal degree matrix (i.e. $D_{ii}=\sum_{j=1}^{n}W_{ij}$). The Unnormalized Graph Laplacian is defined as L=DWL=D-W. Other normalized Laplacian versions are the Symmetric Laplacian $L_{\text{sym}}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ and the Random-Walk (RW) Laplacian $L_{\text{rw}}=D^{-1}L$. Sep-SpectralNet is applicable to all of these Laplacian versions. The eigenvalues of $L$ can be sorted to satisfy $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ with corresponding eigenvectors $v_{1},\dots,v_{n}$ (Von Luxburg, 2007). It is important to note that the first pair (i.e., $\lambda_{1},v_{1}$) is trivial - for every Laplacian matrix $\lambda_{1}=0$, and for the unnormalized and RW Laplacians $v_{1}=\frac{1}{\sqrt{n}}\vec{1}$, namely the constant vector.
@@ 3.2 SpectralNet
The Rayleigh quotient (RQ) of a Laplacian matrix $L\in\mathbb{R}^{n\times n}$ is a function $R_{L}:\mathbb{R}^{n\times k}\to\mathbb{R}$ defined on $A\in\mathbb{R}^{n\times k}$ by
$R_{L}(A)=\operatorname{Tr}(A^{T}LA),$
@@
$\mathcal{L}_{\text{spectralnet}}(\theta)=\frac{1}{m^{2}}R_{L}\big{(}f_{\theta}(X)\big{)}\quad s.t.\quad\frac{1}{m}f_{\theta}(X)^{T}f_{\theta}(X)=I_{k\times k},$ (1)
where $m$ is the minibatch size and $X\in\mathbb{R}^{m\times d}$ contains the $d$-dimensional data samples as rows. Thereby, it learns a map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ (where $d$ is the input dimension) which approximates the space spanned by the first $k$ eigenfunctions of the Laplace-Beltrami operator on the underlying manifold $\mathcal{D}$ (Belkin & Niyogi, 2006; Shi, 2015). Then, it clusters the representations via $k$-means. These eigenfunctions are a natural generalization of the SE to unseen points, enabling both scalable and generalizable spectral clustering.
@@ 3.3 UMAP and Parametric UMAP
Figure 1: Spectral Embedding is crucial for global structure preservation. (a) Non-parametric UMAP preserves the global structure when initialized with SE, but fails to do so with PCA initialization - note the in-cluster separation in the middle row; (b) A similar effect is observed in generalizable implementations: P. UMAP, which does not involve SE, , fails to maintain global structure, resulting in overlapping clusters (e.g., first and second rows). In contrast, NUMAP (ours) achieves global structure preservation comparable to UMAP while supporting generalizability.
UMAP (McInnes et al., 2018) is a widely used visualization method, known for its scalability and its ability to preserve global structure. It constructs a graph from high-dimensional data and learns a low-dimensional representation by minimizing the KL-divergence between the input graph and the representation graph. The UMAP method can be divided into three components (summarized in Fig. 3): (1) constructing a graph which best captures the global structure of the input data; (2) initializing the representations via SE; (3) Learning the representations, via SGD, which best capture the original graph. This setup does not facilitate generalization, as both steps (2) and (3) lack generalizability.
Additionally, as discussed in (Kobak & Linderman, 2021), UMAP primarily derives its global preservation abilities, as well as its consistency, from initializing the representations using SE. Therefore, the SE initialization serves as a critical step for UMAP to uphold the global structure (see demonstrations in Fig. 1a). Global preservation, in this context, refers to the separation of different classes, and avoiding the separation of existing classes. We refer the reader to (Kobak & Linderman, 2021) for a more comprehensive discussion about the effects of informative initialization on UMAP’s performance.
Recently, a generalizable version of UMAP, known as Parametric UMAP (P. UMAP), was introduced (Sainburg et al., 2021). P. UMAP replaces step (3) with the training of a neural network. Importantly, it overlooks step (2), the SE initialization. Consequently, P. UMAP may struggle to preserve global structure, particularly when dealing with non-linear structures. Fig. 1b illustrates this phenomenon using several simple yet non-linear structures. These examples are particularly insightful for highlighting the importance of SE in preserving global structure, as the expected outcome of a good visualization is known. Noticeably, P. UMAP fails to preserve global structure (e.g., it does not separate different clusters).